# Towards AI-Safety-by-Design: A Taxonomy of Runtime Guardrails in Foundation Model based Systems

### v1

@@ Abstract
The rapid advancement and widespread deployment of foundation model (FM) based systems have revolutionized numerous applications across various domains. However, the fast-growing capabilities and autonomy have also raised significant concerns about responsible AI and AI safety. Recently, there have been increasing attention toward implementing guardrails to ensure the runtime behavior of FM-based systems is safe and responsible. Given the early stage of FMs and their applications (such as agents), the design of guardrails have not yet been systematically studied. It remains underexplored which software qualities should be considered when designing guardrails and how these qualities can be ensured from a software architecture perspective. Therefore, in this paper, we present a taxonomy for guardrails to classify and compare the characteristics and design options of guardrails. Our taxonomy is organized into three main categories: the motivation behind adopting runtime guardrails, the quality attributes to consider, and the design options available. This taxonomy provides structured and concrete guidance for making architectural design decisions when designing guardrails and highlights trade-offs arising from the design decisions.
@@
Foundation Model, Large Language Models (LLM), Agent, Guardrails, Safeguard, Responsible AI, AI Safety, Software Architecture, Taxonomy
@@ I Introduction
A foundation model (FM) is a large-scale machine learning model pre-trained on massive amounts of datasets using self-supervision at scale. These models are designed to be highly versatile and can adapt to a wide range of downstream tasks. The term ‘foundation’ reflects their role as the fundamental base upon which many specialized models are built [1]. FMs possess emergent capabilities, meaning their behaviors and susceptibilities arise implicitly from the training data rather than being explicitly programmed. This allows them to perform tasks that were not originally anticipated during their initial training [2].
An FM-based system refers to an application system that utilizes an FM as a core component, interacting with other AI or non-AI components to perform tasks [3]. FM-based systems have experienced significant growth in recent years. For example, approximately 67% of organizations globally have adopted FM-based products [4]. Additionally, their market size is expected to grow at a compound annual growth rate of 36.5% over the next six years [5, 6, 7].
@@
The fast-growing capabilities and autonomy of FM-based systems introduce significant concerns about responsible AI and AI safety, such as generating harmful or offensive content, producing dangerous or unintended outcomes, exhibiting discriminatory behavior, compromising user privacy, spreading misinformation, facilitating cyberattacks, etc [1, 6, 8].
To address these challenges, runtime guardrails are required to be incorporated into FM-based systems to ensure the behavior of the system is responsible and safe. There have been some initial efforts on guardrails, such as input filtering [9, 1], output modification [10, 11], adaptive fail-safes that can prevent harmful outputs [12, 13], real-time monitoring and detection [14, 15, 16, 17], continuous output validation to ensure adherence to ethical standards [18, 19, 20] etc.
@@
However, the existing runtime guardrails mainly focus on inputs and outputs of foundation models, which do not capture the complexity of FM-based systems. FM-based systems are compound AI systems [21], which include foundation models, narrow models for specific tasks, and non-AI components (such as data retriever and external tools). Guardrails need to be placed at different components, taking into account the operational and environmental context. Also, most of the current work primarily addresses functional correctness of guardrails. Quality attributes, such as customizability and interpretability, are often not considered when designing guardrails.
@@
Therefore, in this paper, we present a taxonomy for guardrails from a software architecture perspective. The taxonomy classifies and compares the characteristics and design options of guardrails. There are three main categories in the taxonomy: the motivation for adopting runtime guardrails, the quality attributes that should be addressed in the guardrails design, and the design options available. We developed this taxonomy based on the results of a systematic literature review (SLR). The main contributions of this paper are as follows:
@@
  • The taxonomy offers a comprehensive comparative analysis of different design options and provides structured and concrete guidance for making architectural design decisions about guardrails to achieve AI-safety-by-design. The design options are classified into actions, targets, scopes, rules, autonomy, modalities, and underlying techniques.
@@
  • By summarizing the motivations for adopting runtime guardrails, this taxonomy helps researchers and practitioners understand the underlying rationales and benefits, thereby promoting informed decision-making when integrating guardrails into FM-based systems.
@@
  • Identifying critical quality attributes ensures that the guardrails not only perform their intended functions but also enhance the overall quality of the guardrails’ design.
@@
The rest of the paper is organized as follows: section II discusses background and related works. section III introduces the methodology of this study. The taxonomy of guardrails is presented in section IV answering three key research questions. section V summarizes primary threats to this study. Finally, section VI concludes the paper and outlines the future work.
@@ II Background and Related Work
FMs have significantly impacted the architecture of AI systems [1, 20]. Although guardrails for FM-based systems have been explored in various contexts, this paper is the first study assessing existing work to provide a comprehensive understanding. Below, we present key background and related works required to understand guardrails and the research gap.
@@ II-A Background
FM-based systems incorporate FMs as core components within larger architectures designed to utilize their capabilities [3]. These systems integrate FMs with additional components such as data retrievers (e.g., RAG), external tools and others. The complexity of FM-based systems lies in ensuring seamless interoperability among these components while ensuring responsible AI and AI safety [1].
In 2021, Bommasani et al.[1] provided a comprehensive discussion on foundation models, illustrating the essential elements and relationships between them. The authors also discussed the practical applications of FM-based systems, highlighting their economic and operational advantages over traditional systems. Two years later, Zhou et al. covered recent research advancements, challenges, and opportunities for pre-trained foundation models in text, image, graph, and other data modalities in their survey paper[20]. Both works offer excellent insights into future research directions to address open problems and associated risks. Studies [1, 26] mention that foundation model-based systems often inherit responsible AI and AI safety issues and can cause negative consequences.
Lu et al. developed a taxonomy of FM-based systems focusing on their pretraining, adaptation, architectural design, and responsible-AI-by-design [27]. The taxonomy aids software architects and developers in evaluating and integrating foundation models into complex systems. The authors also highlighted considerations for cost, accuracy, and responsible AI related quality attributes for the systems. The authors later proposed a reference architecture for designing responsible and safe foundation model based systems in [3]. The works [28, 29] explore the risks associated with deploying LLMs and evaluate current approaches to mitigate such risks through model alignment. Recently, Lu et al. discussed that responsible AI practices are highly recommended for FM-based systems [30, 2]. This includes technical solutions and adherence to regulatory and ethical guidelines to maintain public trust and safety [31]. They also emphasized that FM-based systems must ensure they are developed and deployed in a responsible, safe, and legal way [32, 33, 34].
@@ II-B Related Work
There exist several frameworks and tools for designing guardrails at design time and runtime. Significant have been put into model alignment during design time, which focuses on aligning the FM’s outputs with human values and ethical standards. Pre-training and adaptation strategies play a significant role in mitigating risks in FM-based systems through guardrails. Pre-training models on diverse and ethically sourced datasets helps align them more closely with ethical standards. Continuous adaptation through fine-tuning further reduces AI risks [1]. Ensuring ethical and responsible AI practices involves not only technical solutions but also adherence to regulatory and ethical guidelines within the systems. The adoption of comprehensive AI frameworks that incorporate ethical principles into the design and deployment of FMs is crucial for maintaining public trust and safety [31].
Some initial efforts have been made toward runtime guardrails. Zhou et al. [20] present an excellent discussion of the literature that addresses the monitoring and controlling behavior of FM-based systems. Various tools and frameworks have also been developed to tackle these challenges. NeMo Guardrails[16] provides programmable guardrails to ensure that models operate within safe parameters by monitoring inputs and outputs in real-time, dynamically adjusting behavior to prevent harmful outcomes[27]. OpenAI’s Moderation API [35] monitors and filters harmful content generated by models to protect user interactions. The GuardAgent framework [36] utilizes an LLM agent to oversee and safeguard other agents. It demonstrates strong generalization and low operational overhead by dynamically generating guardrail code.
We also found that the continuous validation ensures that the outputs from FM-based systems adhere to predefined ethical standards and guidelines. Techniques such as auditing models [18] through multi-layered approaches [18, 37] systematically check for biases. They also ensure ethical compliance during the model operation [3]. Dynamic access controls adjust access permissions in real time based on the context of data usage to protect sensitive information and ensure it is accessible under appropriate circumstances [38]. Adaptive fail-safes as characteristics of guardrails intervene automatically when an FM exhibits potentially harmful behavior within systems. These fail-safes are designed to modify or halt outputs that could lead to undesirable consequences [31].
However, the design of runtime guardrails for FM-based systems has not been systematically explored. Thus, in this paper, we develop a taxonomy to categorize existing work into a compact framework and compare and evaluate different design options for runtime guardrails.
@@ III Methodology
This study focuses on two primary concepts: (i) foundation model-based systems and (ii) runtime guardrails. We adopted the Petticrew and Roberts approach [39] to define the Population, Interventions, Comparison, Outcomes, and Context (PICOC), within which the intervention in this study is delivered. The PICOC for this study is shown in Table I. Using these PICOC components and following Kitchenham’s guidelines [40], we develop the protocol for this study. The following subsections detail the steps and processes undertaken in this study.
@@
Population Studies and researches focus on runtime guardrails within foundation model-based systems.
Intervention Development, optimization, and evaluation of runtime guardrails in foundation model-based systems, focusing on motivations, key attributes, and design strategies.
Comparison Comparative analysis between different approaches.
Outcomes Design options for runtime guardrails in foundation model based system and their taxonomy structure.
Context Include: Empirical and theoretical studies on the components, design and evaluation of guardrails in foundation model-based systems. Exclude: Studies beyond the scope of foundation model based systems, non-English literature, and those not considering guardrails.
@@ III-A Research Scope and Protocol Development
The high-level research approach for this study is shown in Figure 1. Initially, we determined the research scope and developed a protocol following Kitchenham’s guidelines [40, 41]. The protocol guided the entire study [^1] The protocol and sample data extraction sheet for this study are attached in the supplementary materials by defining relevant scientific databases and resources, formulating keywords and search strings, outlining qualitative and quantitative checklists, and specifying criteria for the inclusion and exclusion of studies.
@@ III-B Research Questions
RQ1: Why are runtime guardrails adopted?
Our first research question investigates the fundamental motivations for designing runtime guardrails in FM-based systems. It explores the responsible AI and AI safety concerns that require the adoption of runtime guardrails to prevent harmful content and unintended behaviors.
RQ2: What are the key quality attributes for designing runtime guardrails?
Our second research question identifies quality attributes that should be considered when designing effective guardrails in FM-based systems.
@@
This question examines the design options for runtime guardrails in FM-based systems from different perspectives, such as action and scope.
@@
Concept 1 (Co1): Foundation Model Based Systems Foundation Models, Foundation Model Based Systems, Large Language Model, Generative AI, Artificial General Intelligence, Transformer Models, Self-supervised Learning, Pretrained Models, Language Models, Conversational AI.
@@ III-C Search string formulation
IC1 Full text of conference papers, journal articles, industry reports, and book chapters that are relevant to the defined main concepts: Foundation model based systems and guardrails.
@@
IC3 Studies that specifically address the design and development of guardrails in foundation model-based systems. This includes theoretical frameworks, empirical research and case studies.
IC4 Papers available in an electronic format, such as PDF, DOC, DOCX, HTML, or PS.
@@
EC2 Discussion papers and opinion pieces that do not provide empirical evidence or concrete solutions related to guardrails in foundation model-based systems.
@@ III-F Data Extraction and Quality Assessment
  • Relevance to guardrails for FM-based systems.
@@
  • Adequate data collection, analysis, and evaluation.
  • Discussion of challenges in designing guardrails.
  • Practical applicability of findings for guardrails.
IV Design Taxonomy for Guardrails in Foundation Model-Based Systems
In this section, we present our taxonomy of runtime guardrails for FM-based systems. As illustrated in Figure 3, this taxonomy is organized into categories based on identified problems, quality attributes, and available design options, which we use in the following sub-sections to answer our key research questions
IV-A Motivation for adopting runtime guardrails (RQ1)
Our first research question investigates the fundamental reasons for integrating guardrails into FM-based systems. We found that guardrails are primarily implemented to mitigate bias, harmful content, and unintended behaviors. Additionally, guardrails better maintain ethical and safety standards in areas such as accuracy, privacy, security, safety, fairness, IP protection, and compliance, as discussed below.
Figure 3: Taxonomy of Guardrails for FM-Based Systems
@@ IV-A1 Accuracy
Accuracy in FM-based systems is a critical concern, particularly regarding the generation of hallucinations, misinformation, and disinformation [45]. Hallucinations occur when models generate information that is factually incorrect or fabricated. Such inaccuracies can mislead users and damage the credibility of the systems [10]. Misinformation refers to the unintentional spread of false information, while disinformation involves the deliberate dissemination of falsehoods to deceive users [2]. These phenomena pose significant risks, as they can propagate false narratives and influence public opinion [20]. Guardrails mitigate these risks by detecting and filtering potentially inaccurate information [16]. For example, OpenAI uses guardrails to clearly label AI-generated content to prevent deepfakes and misinformation [46], such as in one case where it was reported to prevent misleading voters in the upcoming US elections [47, 48]. Therefore, guardrails improve the overall integrity of the systems [46, 1].

### v2

# A Taxonomy of Multi-Layered Runtime Guardrails for Designing Foundation Model-Based Agents: Swiss Cheese Model for AI Safety by Design

@@ Abstract
Foundation Model (FM)-based agents are revolutionizing application development across various domains. However, their rapidly growing capabilities and autonomy have raised significant concerns about AI safety. Designing effective guardrails for these agents is challenging due to their autonomous and non-deterministic behavior, and the involvement of multiple artifacts — such as goals, prompts, plans, tools, knowledge bases, and intermediate and final results. Addressing these unique challenges runtime requires multi-layered guardrails that operate effectively at various levels of the agent architecture, similar to the Swiss Cheese Model. In this paper, we present a taxonomy of multi-layered runtime guardrails to classify and compare their characteristics and design options, grounded on a systematic literature review and guided by the Swiss Cheese Model. This taxonomy is organized into external and internal quality attributes and design options categories. We also highlight the relationships between guardrails, the associated risks they mitigate, and the quality attributes they impact in agent architectures. Thus, the proposed taxonomy provides structured and concrete guidance for making architectural design decisions when implementing multi-layered guardrails while emphasizing the trade-offs inherent in these decisions.
@@
Foundation Model, Large Language Models (LLM), Agent, Guardrails, Swiss Cheese Model, Responsible AI, AI Safety, Software Architecture, Taxonomy
@@ I Introduction
A foundation model (FM) is a large-scale machine learning model pre-trained on massive amounts of data using self-supervision at scale. These models are designed to be highly versatile and can adapt to a wide range of downstream tasks. The term ‘foundation’ reflects their role as the fundamental base upon which many specialized models are built [1]. FMs possess emergent capabilities, meaning their behaviors and susceptibilities arise implicitly from the training data rather than being explicitly programmed. This allows them to perform tasks that were not originally anticipated during their initial training [2]. An FM-based agent utilizes an FM as a core component, interacting with other AI or non-AI components to perform tasks autonomously [3]. In recent years, FM-based agents have experienced extensive growth [4].
@@
The autonomy of FM-based agents along with their enhance decision-making and task execution capabilities introduce significant concerns about responsible AI and AI safety, e.g., generating harmful or offensive content, producing dangerous or unintended outcomes, exhibiting discriminatory behavior, compromising user privacy, spreading misinformation, facilitating cyberattacks, etc [1, 6, 8]. To address these challenges, effective runtime guardrails are required to ensure the agent’s behavior is responsible and safe. Multi-layered runtime guardrails structured similarly to the Swiss Cheese Model that operate at various levels of the agent architecture further mitigate associated risks [76, 1].
@@
There have been some initial efforts on guardrails, such as input filtering [9, 1], output modification [10, 11], adaptive fail-safes that can prevent harmful outputs [12, 13], real-time monitoring and detection [14, 15, 16, 17], continuous output validation to ensure adherence to ethical standards [18, 19, 20] etc. However, the existing runtime guardrails mainly focus on inputs and outputs of foundation models, which do not capture the complexity of FM-based agents. FM-based agents are complex AI systems [21] that integrate foundation models and non-AI components such as data retrievers and external tools. Guardrails need to be placed across different components, taking into account the operational and environmental context, much like the layers in the Swiss cheese model [75]. However, most current work primarily addresses the functional correctness of guardrails, and quality attributes such as customizability and interpretability are often not considered.
@@
Therefore, in this paper, we present a taxonomy for multilayer guardrails from a software architecture perspective. The taxonomy classifies and compares the characteristics and design options of guardrails. There are three main categories in the taxonomy: the external qualities for adopting multi-layered runtime guardrails, the internal attributes that should be addressed in the guardrails design, and the design options available. We developed this taxonomy based on the results of a systematic literature review (SLR), guided by the Swiss Cheese Model structure for AI safety. The main contributions of this paper are as follows:
@@
  • The taxonomy offers a comprehensive comparative analysis of different design options and provides structured and concrete guidance for making architectural design decisions about multi-layered runtime guardrails to achieve AI-safety-by-design. The design options are classified into actions, targets, scopes, rules, autonomy, modalities, and underlying techniques.
@@
  • This taxonomy helps researchers and practitioners understand the underlying rationales and benefits by summarizing the external qualities for multi-layered runtime guardrails. Hence, it also promotes informed decision-making when integrating guardrails into FM-based agents.
@@
  • The taxonomy identify critical internal quality attributes to ensure that the multi-layered runtime guardrails not only perform their intended functions but also enhance the overall quality of the guardrails’ design.
@@
The rest of the paper is organized as follows: section II discusses background and related works. section III introduces the methodology of this study. The taxonomy of guardrails is presented in section IV answering three key research questions. section V presents the implications of our proposed work and summarizes primary threats to this study. Finally, section VI concludes the paper and outlines the future work.
@@ II Background and Related Work
Foundation models (FMs) have significantly impacted the architecture of AI agents [1, 20]. Although guardrails for FM-based agents have been explored in various contexts, this paper is the first study assessing existing work to provide a comprehensive understanding. Below, we present key background and related works required to understand guardrails and the research gap.
@@ II-A Background
FM-based agents incorporate FMs as core components within larger architectures designed to utilize their capabilities [3]. These agents integrate FMs with additional components such as data retrievers (e.g., RAG), external tools and others. The complexity of FM-based agents lies in ensuring seamless interoperability among these components while ensuring responsible AI and AI safety [1].
In 2021, Bommasani et al.[1] provided a comprehensive discussion on foundation models, illustrating the essential elements and relationships between them. The authors also discussed the practical applications of FM-based agents, highlighting their economic and operational advantages over traditional systems. Two years later, Zhou et al. covered recent research advancements, challenges, and opportunities for pre-trained foundation models in text, image, graph, and other data modalities in their survey paper[20]. Both works offer excellent insights into future research directions to address open problems and associated risks. Studies [1, 26] mention that foundation model-based agents often inherit responsible AI and AI safety issues and can cause negative consequences.
Lu et al. developed a taxonomy of FM-based systems (agents) focusing on their pretraining, adaptation, architectural design, and responsible-AI-by-design [27]. The taxonomy aids software architects and developers in evaluating and integrating foundation models into complex systems. The authors also highlighted considerations for cost, accuracy, and responsible AI related quality attributes for the systems. The authors later proposed a reference architecture for designing responsible and safe foundation model based systems (agents) in [3]. The works [28, 29] explore the risks associated with deploying LLMs and evaluate current approaches to mitigate such risks through model alignment. Recently, Lu et al. discussed that responsible AI practices are highly recommended for FM-based systems (agents) [30, 2]. They also emphasized that FM-based agents must ensure they are developed and deployed in a responsible, safe, and legal way [32, 33, 34, 31].
@@ II-B Related Work
There exist several frameworks and tools for designing guardrails at design time and runtime. Significant effort have been put into model alignment during design time, which focuses on aligning the FM’s outputs with human values and ethical standards. Pre-training and adaptation strategies play a significant role in mitigating risks in FM-based agents through guardrails, especially multilayer guardrails. Pre-training models on diverse and ethically sourced datasets helps align them more closely with ethical standards. Continuous adaptation through fine-tuning further reduces AI risks [1]. Ensuring ethical and responsible AI practices involves not only technical solutions but also adherence to regulatory and ethical guidelines within the systems. The adoption of comprehensive AI frameworks that incorporate ethical principles into the design and deployment of FMs is crucial for maintaining public trust and safety [31].
Some initial efforts have been made toward runtime guardrails. Zhou et al. [20] present an excellent discussion of the literature that addresses the monitoring and controlling behavior of FM-based agents. Various tools and frameworks have also been developed to tackle these challenges. NeMo Guardrails[16] provides programmable guardrails to ensure that models operate within safe parameters by monitoring inputs and outputs in real-time, dynamically adjusting behavior to prevent harmful outcomes[27]. OpenAI’s Moderation API [35] monitors and filters harmful content generated by models to protect user interactions. The GuardAgent framework [36] utilizes an LLM agent to oversee and safeguard other agents. It demonstrates strong generalization and low operational overhead by dynamically generating guardrail code.
We also found that the continuous validation ensures that the outputs from FM-based agents adhere to predefined ethical standards and guidelines. Techniques such as auditing models [18] through multi-layered approaches [18, 37] systematically check for biases. They also ensure ethical compliance during the agent’s operation [3]. Dynamic access controls adjust access permissions in real time based on the context of data usage to protect sensitive information and ensure it is accessible under appropriate circumstances [38]. Adaptive fail-safes as characteristics of guardrails intervene automatically when an FM exhibits potentially harmful behavior within agents. These fail-safes are designed to modify or halt outputs that could lead to undesirable consequences [31]. However, the design of multi-layered runtime guardrails for FM-based agents has not been systematically explored. Thus, we develop a taxonomy to categorize existing work into a compact framework and compare and evaluate different design options for multi-layered runtime guardrails.
@@ III Methodology
This study focuses on two primary concepts: (i) foundation model-based agents and (ii) runtime guardrails. We adopted the Petticrew and Roberts approach [39] to define the Population, Interventions, Comparison, Outcomes, and Context (PICOC), within which the intervention in this study is delivered. The PICOC for this study is shown in Table I. Using these PICOC components and following Kitchenham’s guidelines [40], we develop the protocol for this study. The following subsections detail the steps and processes undertaken in this study.
@@
Population Studies and researches focus on runtime guardrails within foundation model-based agents.
Intervention Development, optimization, and evaluation of multilayer runtime guardrails in foundation model-based agents, focusing on key quality attributes and design strategies inspired by the Swiss Cheese Model.
Comparison Comparative analysis between different approaches to design multilayer runtime guardrails in FM-based agents.
Outcomes Qualities and design options for multilayer runtime guardrails for foundation model-based agents guided by Swiss cheese model.
Context Include: Empirical and theoretical studies on the components, design and evaluation of guardrails in foundation model-based agents. Exclude: Studies beyond the scope of foundation model based agents, non-English literature, and those not considering guardrails.
@@ III-A Research Scope and Protocol Development
The high-level research approach for this study is shown in Figure 1. Initially, we determined the research scope and developed a protocol following Kitchenham’s guidelines [40, 41]. The protocol guided the entire study by defining relevant scientific databases and resources, formulating keywords and search strings, outlining qualitative and quantitative checklists, and specifying criteria for the inclusion and exclusion of studies.
@@ III-B Research Questions
RQ1: External qualities for runtime multilayer guardrails in FM-based agents?
Our first research question investigates the fundamental external qualities for designing multilayer runtime guardrails in FM-based agents. It explores the responsible AI and AI safety concerns that require the adoption of runtime guardrails to prevent harmful content and unintended behaviors.
RQ2: What are the key internal quality attributes for designing multilayer runtime guardrails?
Our second research question identifies quality attributes that should be considered when designing effective guardrails in FM-based agents.
@@
This question examines the design options for runtime guardrails in FM-based agents from different perspectives, such as action and scope, guided by the Swiss Cheese Model.
@@
Concept 1 (Co1): Foundation Model based agents Foundation Models, Foundation Model based agents, Large Language Model, Generative AI, Artificial General Intelligence, Transformer Models, Self-supervised Learning, Pretrained Models, Language Models, Conversational AI.
@@ III-C Search String Formulation
IC1 Full text of conference papers, journal articles, industry reports, and book chapters that are relevant to the defined main concepts: Foundation model based agents and guardrails.
@@
IC3 Studies that specifically address the design and development of guardrails in foundation model-based agents. This includes theoretical frameworks, empirical research and case studies.
IC4 Papers available in an electronic format, such as PDF, DOC, DOCX, HTML, and PS etc.
@@
EC2 Discussion papers and opinion pieces that do not provide empirical evidence or concrete solutions related to guardrails in foundation model-based agents.
@@ III-F Data Extraction and Quality Assessment
  • Relevance to guardrails for FM-based agents.
@@
  • Adequate data collection, analysis, and evaluation of guardrail effectiveness at different layers of the agent architecture.
  • Discussion of challenges in designing guardrails for autonomous and non-deterministic behaviors in agents.
  • Practical applicability of findings for guardrails in FM-based agents.
IV The Proposed Taxonomy
In this section, we present our taxonomy of multilayer runtime guardrails for FM-based agents. As illustrated in Figure 3, this taxonomy is organized into three main categories: external quality attributes, internal quality attributes, and design options, which we use in the following sub-sections to answer our key research questions guided by the Swiss Cheese Model for AI Safety by Design.
IV-A External Qualities for Multi-layered Runtime Guardrails (RQ1)
Our first research question investigates the fundamental reasons for integrating guardrails into FM-based agents. We found that guardrails are primarily implemented to mitigate bias, harmful content, and unintended behaviors. Additionally, guardrails better maintain ethical and safety standards in areas such as accuracy, privacy, security, safety, fairness, cost, performance and IP protection, as discussed below.
Figure 3: Taxonomy of Multi-Layered Runtime Guardrails for FM-Based Agents
@@ IV-A1 Accuracy
Accuracy in FM-based agents is a critical concern, particularly regarding the generation of hallucinations, misinformation, and disinformation [45]. Hallucinations occur when models generate information that is factually incorrect or fabricated. Such inaccuracies can mislead users and damage the credibility of the systems [10]. Misinformation refers to the unintentional spread of false information, while disinformation involves the deliberate dissemination of falsehoods to deceive users [2]. These phenomena pose significant risks, as they can propagate false narratives and influence public opinion [20]. Guardrails mitigate these risks by detecting and filtering potentially inaccurate information [16]. For example, OpenAI uses guardrails to clearly label AI-generated content to prevent deepfakes and misinformation [46], such as in one case where it was reported to prevent misleading voters in the upcoming US elections [47, 48]. Therefore, guardrails improve the overall integrity of the systems [46, 1].